OpenStudy (ragingsquirrel):

If the probability of an event is 26%, what is the probability of its complement occurring? 100% 74% 26% 24% Please explain compliments.

OpenStudy (ash2326):

P(happening)+P(not happening)=1 P(happening)=26% or 0.26 P(not happening)=1-0.26=0.74 so 74%

OpenStudy (ragingsquirrel):

So the compliment is just the chance of it not happening?

OpenStudy (ash2326):

yeah

OpenStudy (ragingsquirrel):

Thanks!